# Confinement-induced collective motion in suspensions of run-and-tumble particles

### v1

@@ III.4 Stability of the band with respect to $\theta_{obs}$
Interestingly, for $22.5^{\circ}\leq\theta_{obs}\leq\ 30^{\circ}$ the traveling band is fully formed and reaches its top speed. As we highlighted before, the counter-current of active particles in domain III is the facilitator of the the collective motion, and thus, when we monitor $\Delta_{x}(y)$ as $\theta_{obs}$ varies, we observe a non-monotonic behavior that perfectly correlates with the behavior of $\langle V_{CM}\rangle$, exhibiting the highest peaks, precisely, for $\theta_{obs}=22.5^{\circ}$ and $30^{\circ}$. Thus, the larger the counter-current the faster the traveling band, which might seem counter-intuitive in the first place. However, since $\Delta_{x}(y)$ accounts for rearrangement and diffusive fluxes opposing to the band’s motion, we suggest that those angles favor particles’ rearrangements at the lower part of the band by enhancing sliding and thus producing more vacancies for particles in domain III to diffuse into.
@@ IV Summary and Conclusions
In the reported work, we have studied the collective behavior of persistent active suspensions confined in asymmetric channels. The top boundary of the channel is a flat wall and its bottom boundary is composed of an array of funnel-like obstacles (see Fig. 1). The active particles follow translational overdamped Langevin’s dynamics and rotational run-and-tumble dynamics. Our numerical results show a new type of collective motion in the shape of a wide and dense traveling band induced by the symmetry-breaking of funnel-like obstacles. Such traveling band has not been observed neither in bulk nor under confinement in channels bounded by two flat walls.
@@
Interestingly, we found the existence of a narrow range of obstacles tilt angles, $\theta_{obs}$ which promotes the emergence of the traveling state. We discuss the preferred $\theta_{obs}$ in terms of an enhanced counter-current, $\Delta_{x}(y)$, as a consequence of an improved sliding along the obstacles for such angles.
We suggest that a possible experimental realization of the system might consist in a concentrated suspension of the motile microalgae Chlamydomonas reinhardtii confined into a ring-shaped microfluidic device with one asymmetric boundary (with the shape of corrugations or funnels). It has been shown that a single cell of C. reinhardtii persistently follows the curvature of a circular chamberbentley2022phenotyping, which suggests that a collection of cells might behave in a similar way as described in the article. The ring-shaped chamber should have a perimeter of the same order of magnitude of the persistence length of the microalgae. To achieve high persistence lengths and concentrate the cells in specific regions within the ring, C. reinhardtii phototaxis can be leveraged. Another possible experimental realization of the proposed mechanism consists in Janus particles confined into rings. A recent study has reported clustering of Janus particles in ring-shaped soft confinement knippenberg2024motility. Such clustering effect might be the basis for a traveling structure when the width of the ring and the packing fraction of the active suspension are correctly tuned. A potential technological application of the confinement-induced tracked locomotion could be fast steered one-dimensional transport of passive particles in narrow channels.
@@ V Supplementary Information
  • SI containing additional figures and discussions. Figure S1: bulk, channels with 2 flat walls and channels with corrugations with no gaps. Fig. S2: pair correlation functions to choose the correct cut-off radius for the percolation criterion. Figure S3: Finite size effects, proof that in long channels the traveling band is still formed.
@@ Acknowledgements.
C.V. acknowledges fundings IHRC22/00002 and Proyecto PID2022-140407NB-C21 funded by MCIN/AEI /10.13039/501100011033 and FEDER, UE. C.V. and H.S. acknowledge fundings from the European Union’s Horizon research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101108868 (BIOMICAR). H.S and J. M-R. thanks the insightful discussions with Rodrigo Fernández-Quevedo García. H. S. and C. V. acknowledge the discussions with Sujeet Kumar Choudhary and Marco Polin.

### v2

@@ III.4 Stability of the band with respect to confining geometry
Let us start discussing the influence of $\theta_{obs}$. We perform simulations at $\phi=0.30$ and $l_{p}=10^{5}$ for $\theta_{obs}\in[0^{o},75^{o}]$. In Fig. 6a, using $\langle V_{CM}\rangle$ as an order parameter, we observe the existence of a range $15^{o}\leq\theta_{obs}\leq 45^{o}$ which allows the formation of traveling structures. Interestingly, for $22.5^{\circ}\leq\theta_{obs}\leq\ 30^{\circ}$ the traveling band is fully formed and reaches its top speed. As we highlighted before, the counter-current of active particles in domain III is the facilitator of the the collective motion, and thus, when we monitor $\Delta_{x}(y)$ as $\theta_{obs}$ varies, we observe a non-monotonic behavior that perfectly correlates with the behavior of $\langle V_{CM}\rangle$, exhibiting the highest peaks, precisely, for $\theta_{obs}=22.5^{\circ}$ and $30^{\circ}$. Thus, the larger the counter-current the faster the traveling band, which might seem counter-intuitive in the first place. However, since $\Delta_{x}(y)$ accounts for rearrangement and diffusive fluxes opposing to the band’s motion, we suggest that those angles favor particles’ rearrangements at the lower part of the band by enhancing sliding and thus producing more vacancies for particles in domain III to diffuse into. We also consider channels in which the period of the obstacles at the bottom wall, $l_{obs}$, is increased. To do so, we need to vary both $L_{obs}$ and $\theta_{obs}$ simultaneously to keep the amplitude of the obstacles, $A=L_{obs}\sin(\theta_{obs}+\sigma_{22})/2$, constant. We take the case used to built the phase diagram in Fig. 2a as reference, with $\theta_{obs}=30^{o}$, $L_{obs}=3\sigma_{22}$ and $d=1.25\sigma_{22}$, for which $l^{ref}_{obs}=9.70\sigma_{22}$. We study cases with $l_{obs}=nl^{ref}_{obs}$, $n$ being a positive integer, and packing fraction, $\phi=0.30$. We observe that increasing $l_{obs}$ in general favors clogging. At $l_{obs}=2l_{obs}^{ref}$, a traveling band appears for $2300\tau$ to finally form a clog. Note that as $l_{obs}$ increases, the geometry of the channel approaches the limit of two flat walls that, as already discussed, only promotes the formation of clogging states. Examples of these systems can be found in Supplementary Video 5
Next, we consider a special case in which the top wall is gradually deformed by increasing the angle $\beta$ (see Fig. 6b), which converts the top wall into a trapezoidal wave, while keeping $\theta_{obs}=30^{o}$. In the inset of Fig. 6b, in the upper panel, we present the geometry of the channel when $\beta=0^{o}$, whereas, in the bottom panel, we present the geometry of the channel for an arbitrary $\beta>0^{o}$. As for the funnel-like obstacles in the bottom wall for which $y=0$ is placed to cross the mid points of the obstacles, $y=L_{y}$ is placed to cross the mid points of the trapezoids. Fig. 6b shows that confinement-induced tracked locomotion only takes place for very low $\beta$ angles. For $\beta\geq 7.5^{o}$, the active particles self-organize into a big clog. On the one hand, very small values of $\beta$ resembles the case of a flat top wall and thus recovers the traveling structure, remarking the importance of asymmetric walls to stabilize it. On the other hand, as $\beta$ increases, the active particles cannot slide freely along the top wall anymore and experience a higher effective friction coefficient when move close to the top wall due to the presence of the trapezoids, which promotes the formation of clogs. Regarding the stability of the traveling bands, now we can say that the sliding along the top walls of particles belonging to domains I and II is fundamental to sustain the confinement-induced tracked locomotion. The microchannels in the inset are inside colored rectangles following the same color code as in the phase diagram of Fig. 2a (red for traveling states and green fro clogging states), and thus indicating that channels with small $\beta$ favor the formation of the traveling band, whereas higher values favor the formation of clogs.
Finally, we study the influence of the channels width, $L_{y}$ on the formation of traveling bands. We perform simulations of systems with $\theta_{obs}=30^{o}$ and $L_{x}=797\sigma_{22}$ while varying $L_{y}\in[5\sigma_{11},40\sigma_{11}]$. We report $\langle V_{CM}\rangle$ as a function of $L_{y}$ in Fig. 6c. First, we note that very narrow channels only allow the formation of clogs, which is reflected in $\langle V_{CM}\rangle\approx 0$. When $\phi=0.30$, we observed fully-developed confinement-induced tracked locomotion for $10\sigma_{11}\leq L_{y}\leq 15\sigma_{11}$ , whereas for $20\sigma_{11}\leq L_{y}\leq 25\sigma_{11}$ a big cluster is formed and slides along the top wall. Even though the cluster is large, its size is not enough to span the channel in the $y$ direction and thus forming a traveling band. Such a cluster reaches speeds around $10\%$ of the self-propulsion speed of individual particles, and can be understood as a precursor of the traveling band only containing domains I and II (see Fig.4c). For $L_{y}\geq 30\sigma_{11}$ a clustering state with no persistence, as that one described in the phase diagram of Fig. 2a is observed. We have learned that the size of the cluster containing domains I and II is important to the formation of traveling states within the microchannel. In order to get more insights on this matter, we repeat the previous analysis but now for $\phi=0.5$ to induce the formation of larger clusters. As expected, at $\phi=0.50$ those states that correspond to a persistent cluster sliding on the top wall at $\phi=0.30$ become fully-developed traveling structures, as can be seen, for example, for $L_{y}=20\sigma_{11}$. Again, as $L_{y}$ increases, the fully-developed traveling state converts into a persistent cluster to finally become a non-persistent clustering state. Although these are still preliminary results, we can state, to certain extent, that increasing $L_{y}$ shifts the phase behavior of the system towards higher values of $\phi$.
@@ IV Summary and Conclusions
In this article, we have studied the collective behavior of persistent active suspensions confined in asymmetric channels. The top boundary of the channel is a flat wall and its bottom boundary is composed of an array of funnel-like obstacles (see Fig. 1). The active particles follow translational overdamped Langevin’s dynamics and rotational run-and-tumble dynamics. Our numerical results show a new type of collective motion in the shape of a wide and dense traveling band induced by the asymmetry of the boundaries of the confining channel. Such traveling band has not been observed neither in bulk nor under confinement in channels bounded by two flat walls.
@@
Interestingly, we found the existence of a narrow range of obstacles tilt angles, $\theta_{obs}$ which promotes the emergence of the traveling state. We discuss the preferred $\theta_{obs}$ in terms of an enhanced counter-current, $\Delta_{x}(y)$, as a consequence of an improved sliding along the obstacles for such angles. We have discussed the effects of obstacles with a well-defined period on the top wall, by including an angle $\beta$ that gradually converts the top wall from flat to a trapezoidal wave. We observed that only small $\beta$ allows for the formation of the traveling structures, since the sliding of domains I and II along the top wall is crucial to the development of traveling states and high $\beta$ angles hinder the motility of active particles by promoting their accumulation inside the trapezoids, and effectively increasing the friction coefficient when particles slide along the top wall. Finally, we addressed the effects of increasing the channels’ width, $L_{y}$, finding that in wider channels the active particles’ packing fraction must be increased to recover the formation of traveling bands.
We suggest that a possible experimental realization of the system might consist in a concentrated suspension of the motile microalgae Chlamydomonas reinhardtii confined into a ring-shaped microfluidic device with one asymmetric boundary (with the shape of corrugations or funnels). It has been shown that a single cell of C. reinhardtii persistently follows the curvature of a circular chamberBentley et al. (2022), which suggests that a collection of cells might behave in a similar way as described in the article. The ring-shaped chamber should have a perimeter of the same order of magnitude of the persistence length of the microalgae. To achieve high persistence lengths and concentrate the cells in specific regions within the ring, C. reinhardtii phototaxis can be leveraged. Another possible experimental realization of the proposed mechanism consists in Janus particles confined into rings. A recent study has reported clustering of Janus particles in ring-shaped soft confinement Knippenberg et al. (2024). Such clustering effect might be the basis for a traveling structure when the width of the ring and the packing fraction of the active suspension are correctly tuned. The confinement-induced tracked locomotion described in this article is persistent in long enough timescales (see Figure S4) to be appealing for technology. A potential technological application could be the fast steered one-dimensional transport of passive objects along narrow channels. In the coming future, we plan to address a detailed characterization of the transition to clogging states since it is relevant in granular flows in confined environments Arévalo and Zuriguel (2016); Sinha et al. (2025); Fang et al. (2024), and might be of interest to understand biofilms formation under confinement Kumar et al. (2013); Conrad and Poling-Skutvik (2018); Fortune et al. (2022). Another interesting study will be inducing preferential directions on the confinement-induced tracked locomotion by adjusting the shape of the obstacles.
@@ V Supplementary Information
  • SI containing additional figures and discussions. Figure S1: bulk, channels with 2 flat walls and channels with corrugations with no gaps. Fig. S2: pair correlation functions to choose the correct cut-off radius for the percolation criterion. Figure S3: Finite size effects, proof that in long channels the traveling band is still formed. Figure S4: results of the time evolution of the center of mass of the system for a long simulation, showing inversion events and transient clogging states
@@
  • Supplementary Video 4: a movie showing the active suspension confined into asymmetric channels as those presented in Fig. 6b: trapezoids with a characteristic angle $\beta$ in the top wall and funnel-like obstacles with an angle $\theta_{obs}$ in the bottom wall. $\theta_{obs}=30^{o}$ and $\beta=5^{o}$ (top), $30^{o}$ (central) and $60^{o}$ (bottom).
  • Supplementary Video 5: a movie showing the active suspension confined into microchannels in which the period of the obstacles in the bottom wall, $l_{obs}$, is increased. $l_{obs}^{0}=2\left(L_{obs}\cos\theta_{obs}+d+\sigma_{22}\right)$ with $\theta_{obs}=30^{o}$, $L_{obs=3\sigma_{22}}$ and $d=1.25\sigma_{22}$. In the movie, $l_{obs}=2l_{obs}^{0}$ (top), $l_{obs}=3l_{obs}^{0}$ (central) and $l_{obs}=4l_{obs}^{0}$ (bottom)
@@ Acknowledgements.
C.V. acknowledges fundings IHRC22/00002 and Proyecto PID2022-140407NB-C21 funded by MCIN/AEI /10.13039/501100011033 and FEDER, UE. C.V. and H.S. acknowledge fundings from the European Union’s Horizon research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101108868 (BIOMICAR). H.S and J. M-R. thanks the insightful discussions with Rodrigo Fernández-Quevedo García. H. S. and C. V. acknowledge the discussions with Sujeet Kumar Choudhary and Marco Polin. The authors acknowledge insightful suggestions by Hugues Chaté.